The difference in Coulomb energyfor nuclie with Z+1 and Z protons is given by?
1
Expert's answer
2020-06-11T10:35:29-0400

The difference in Coulomb energy for nuclei with Z+1 and Z protons is given by


"\text{d}E_C=\frac{1}{4\pi\epsilon_0}\frac{6}{5}\frac{Ze^2}{R}."